In December of 20x4, Memphis Plastics sold a forklift for $18,300. The machine was purchased in 20x1 for $80,000. Since then2. Compute the gain or loss on the sale on sale3. Determine the after-tax cash flow at the time the forklift was sold. The firms tax rate is 45 percent After-tax cash flow
Answer #1

1.

Book value = purchased cost - accumulated depreciation

= 80,000 - 59,200

= $20,800

2. loss

= sale - book value of machine

= 18,300 - 20,800

= $ -2500 or $(2500)

3.

After tax cash flow = 18,300 - 45%

= $10,065